the and Convolution to find Use Laplace transformt solution to the + % IVP y 24+ 5y = s y lozo, y/o = -2
Answer #1

2/3 t - Given that y - 2y + 5y yo)=0 y(o)= -2 uring loplace tonoforon. Lff (t) = f(s) = f(t) e set det from formula Lyl - Son both side in Now applying Laplace transform eqreation a L[Eg - 29 + sy] = L(42] 8 L[y] - Syo) - gro) SLCYJ + YO) +5l[y]-> 52-5+5) opplying convolution theorem → 1?$£$) esy=[-$4893x C- Sars)3 Now applying Here leat f(3) = 1513 and rs)= 5513 S²S+15/3 2/3 26 1 24 ein (5+) $513 ( 82-5+5) والے Now YA) = 24 24 3 Y sin nig sin ( t ) - 2x2 e ainting at Vagin ze pin (E ) () a

If you have any doubt then please comment.